Cutter Enterprises purchased equipment for $81,000 on January 1, 2018. The equipment is expected to have a five-year life and a residual value of $6,300. Using the double-declining balance method, depreciation for 2018 and the book value at December 31, 2018, would be:
Answer #1
The percentage of depreciation = (100%/5)*2 = 40.00%
Depreciation for 2018, the first year with half year convention = 40%/2 = 20.00%
Depreciation for 2018 = 81000*20% = $               16,200
Book value on December 31, 81000-16200 = $               64,800

With full year convention the figures would be as below:

The percentage of depreciation = (100%/5)*2 = 40.00%
Depreciation for 2018, the first year with full year convention 40.00%
Depreciation for 2018 = 81000*20% = $               32,400
Book value on December 31, 81000-32400 = $               48,600